Adapted from Algebra and Trigonometry by P. Forester but I have a textbook by Holt
Prove the product of two negatives is a positive
(-x)(-y) = (-1*x)(-1*y) Identity
= (-1)[x*(-1)](y) distributive
= (-1)[-1*x](y)
= [-1*(-1)] xy
= 1*xy
= xy
therefore (-x)(-y) = xy

Answer by stanbon(75887)

We know -(xy) + xy = 0 by definition of additive inverse
--------------------------
Now show -(-xy) + xy = 0 where x and y are both positive numbers.
--------------------------
(-x)(-y) = (-1)(x)(-y) -----associative law
(-1)[(x)(-y)]=(-1)(-xy)--associative law & (positive * negative is negative)
(-1)(-xy) + (-xy) = 0 definition of additive inverse
Therefore (-1)(-xy) = xy ... uniqueness of additive inverse
Therefore (-x)(-y) = xy ... argument above
==================
Cheers,
Stan H.